OpenStudy (anonymous):

A 25 ounce solution is 20 percent alcohol,if 50 ounces of water are added to it,what percent of the new solution is alchohol?

OpenStudy (anonymous):

if your 25 ounces is 20% booze, then you have 20% of 25 = \(0.2\times 25=5\) ounces of booze now you add 50 ounces of water, giving you \(25+50=75\) ounces of liquid, of which 5 ounces is booze your job is to figure out what 5 out of 75 is as a percent compute \(\frac{5}{75}\) and then move the decimal two places left